Suppose the economy is operating at equilibrium, with Y0 5 1,000. If the government undertakes a fiscal change whereby the tax rate, t , increases by .05 and government spending increases by 50, will the budget surplus go up or down? Why?


1
Expert's answer
2021-08-20T08:49:59-0400

Given,

Equilibrium income (Y0) = 1,000

Increase in tax rate (t) = 0.05

Increase in government spending = 50


Budget surplus: budget surplus is the difference between government revenue (taxes) and government spending.


Change in Budget surplus=Change in tax revenue−Change in Government spending

Change in tax revenue=t×Y0

Change in tax revenue=0.05×1000

Change in tax revenue=50

Change in Government spending=50


Now,

Change in budget surplus=Change in tax revenue−Change in Government

 spending

Change in budget surplus=50 −50

Change in budget surplus=0

Change in Budget surplus=Change in tax revenue-Change in Government spending Change in tax revenue=t×Y0

Change in tax revenue=0.05×1000Change in tax revenue=50Change in Government spending=50Now,Change in budget surplus=Change in tax revenue-Change in Government spending Change in budget surplus=50 -50Change in budget surplus=0

The change in budget surplus is 0. So the budget surplus will remain unchanged.


Answer: The budget surplus will not change